What is the probability of rolling the sum of 5 with two dices?
With 2 dice, there are 36 outcomes
Of those, a sum of 5 can be made in 4 ways (1 + 4, 2 + 3, 3 + 2, 4 + 1)
⇒ pr(5) = 4/36
= 1/9
≈ 0.11
≈ 11 %
What is the probability that if you draw 5 marbles from the jar without replacement 3 are red?
That depends on how many marbles are in the jar, and what colours they are. You could have 999,997 black marbles and 3 red ones, which would make the chance of drawing one red marble, quite literally, one in one million.
As an interesting aside, as the number of non-red marbles grows, the probability of drawing a red one gets infinitely closer to (but never reaches) zero, assuming there is at least one red marble. This is an example of something from calculus, which becomes quite useful in thinking about some kinds of things.
Please note that WikiAnswers is not here to do your homework for you. If this question is from work you have been assigned by a teacher or professor, I suggest you do the work yourself instead, which will help to increase your understanding of the subject matter and help you to become more familiar with it.
A one-time pad is an encryption/deencryption method where a random series of cipher codes is used one time, and one time only, in a communication. Assuming that only the two parties possess their pairs of pads, then the encryption is unbreakable.
How many 1s will you get if you roll a die 120 times?
If it's a fair cube, then the probability of a ' 1 ' (or any other number 2 thru 6) on
any roll is 1/6 . The most probable result of 120 rolls is 20 1s, 20 2s, 20 3s, 20 4s,
20 5s, and 20 6s.
That's the most probable. This means that if you do 120 rolls an infinite number
of times, that's the set of results you'll get most often. But it does NOT tell you
what you WILL get if you do 120 rolls one time.
A related comment: If you roll the cube 119 times and a ' 1 ' has never come up
yet ... and it's a fair and honest cube ... then the chance of rolling a ' 1 ' on the
120th roll is . . . . (wait for it) . . . . 1/6 , just like it was on the first roll. The cube
has no memory, and there's no such thing as "it's due".
First draw probability of red is 6/10.
If first draw is red, then second draw probability of red is 5/9.
Multiply the two probabilities to get the overall probability.
6/10 X 5/9 = 30/90 = 1/3
Why wood u say that because that .
What lf you rolled a two digit number in probability?
It is no big deal. You could do that if the die was not the standard cube with numbers 1 to 6.
What is the size of a sample for a survey?
The population is every data point you intend to generalise the survey results to. The sample frame is those data points that you can pick from for the survey. The sample is which of these data points you actually survey, and the sample size is how many of those data points there are.
For instance, if you have 700 students in a school, and you have access to 300 of them, and decide to give 30 of them a survey, the sample size is 30.
How long would it take to drive to Orlando Florida from Lynchburg Virginia going 70 miles an hour?
Orlando Florida is about 705 miles from Lynchburg Virginia. At 70 miles per hour, that would take about 10.5 hours.
Mapquest says 11 hours and 40 minutes, but that is accounting for speed limits and routing.
What is the probability of get blue if there 4blue 6yellow 3red?
The probability of blue, when there are 4 blue, 6 yellow, and 3 red, is 4 is 13.
The probability of any event is a number in the interval [0,1 ]. Therefore, in probability, 8 is very a mistake.
That area has few plate forms around it.
What word tells the value found by subtracting the mean and dividing by the standard deviation?
z score
What is the probaaility of drawing a face card that is not a heart?
There are 12 face cards and three of them are hearts. Therefore, the probability of drawing a face card that is heart is 3/12 = 1/4. The probability of drawing a face card that is not a heart is 1-1/4 = 3/4.
How can you make a better Yugioh deck with the 3 Egyptian gods and the 3 wicked cards and dragons?
That would make a bad deck, considering all that would require too many Tributes
What is the probability of rolling 3 straights in a row?
There is insufficient information in the question to answer it. You did not specify how many dice were involved. Please restate the question.
In the meantime, I will give a generalized answer.
If the probability of rolling a straight is n, then the probability of rolling straight three times in a row is n3. For instance, with three dice, the probability of rolling a straight is 1 in 36, neglecting (i.e. allowing for) wrap-around straights, and ignoring (sorry, but the math is too complex and does not serve the answer anyway) the fact that straights can go in either direction. That 1 in 36 is n=0.02778. The probability of doing that 3 times in a row, then, is 0.00002143.